I have two more sisters than brothers.Each of my sisters has two more sisters than brothers.How many more sisters than brothers does my youngest brother have?
Answer by ikleyn(52788)

.
1.  It follows from the first two sentences that the child who says it, is a girl.



2.  So, the girl has two more sisters than brothers:


    g -1 = b + 2,    or   g = b + 3,


    where "g" is the number of girls and "b" is the number of boys in the family.


    Hence, each boy (each brother) has 4 more sisters than brothers.

--------------

Formally, I just answered the question and solved the problem.

But I'd like to highlight that this answer does not pre-define a unique number of girls and boys in the family -
these numbers can vary.

For example, possible configuration might be (5 girls,2 boys) or (6 girls, 3 boys),

but in any case the difference the question asks for is 4.

---------------

Also, as an accompanying fact, notice that the number of boys in the family is "at least 2", from the condition.

It is not required by the problem, but it is useful to derive it and to keep it in your mind in order for the problem makes sense.